[figure: 9.7 cm]

ELECTRONIC STRUCTURE OF CUBIC DyCu_{5}

B. Idzikowski^{a}, G. Chełkowska^{b}, A. Jezierski^{a}

aInstitute of Molecular Physics, Polish Academy of Sciences, 60-179 Poznań, Poland bInstitute of Physics, University of Silesia, 40-007 Katowice, Poland

Binary RCu_{5} (R=heavy rare earth) compounds can crystallize in two different allotropic structures: (i) low-temperature cubic and (ii) high-temperature hexagonal forms. We report on the electronic properties of Dy-Cu compound in cubic AuBe_{5}-type structure. The single phase sample was produced in Ar atmosphere using conventional one wheel melt-spinning equipment as long ribbon. The valence band spectra of the compound were investigated by X-ray photoelectron spectroscopy (XPS) and ultraviolet photoelectron spectroscopy (UPS). Results of ab initio calculation of the electronic structure using the tight-binding linear muffin-tin orbital (TB LMTO) method are compared with experimental data. This work was developed within the Grant No. KBN 2 P03B 115 22 of the Polish Committee for Scientific Research.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Bogdan Idzikowski

Address for correspondence : Institute of Molecular Physics Polish Academy of Sciences M. Smoluchowskiego 17 PL 60-179 Poznań, Poland

Email address : idzi@ifmpan.poznan.pl